# Non-Homophilic Graph Pre-Training and Prompt Learning

### v4

@@ A. Algorithm
We detail the main steps for conditional prompt generation and tuning in Algorithm 1. In brief, we iterate through each downstream task to learn the corresponding prompt vectors individually. In lines 3–5, we compute the embedding for each node using the pre-trained graph encoder, with the pre-trained weights $\Theta_{0}$ remaining fixed throughout the adaptation process. In lines 8–22, we optimize the condition-net. Specifically, we performt similarity-weighted readout (lines 9–11), generate prompts (lines 12–13), modify nodes’ embeddings using these prompts (lines 12–15), and update the embeddings for the prototypical nodes/graphs based on the few-shot labeled data provided in the task (lines 18–19). Note that updating prototypical nodes/graphs is necessary only for classification tasks.
@@ B. Homophily and Non-Homophily Methods
We provide further details about the set of positive samples $\mathcal{A}$ and negative samples $\mathcal{B}$ of homophily and non-homophily methods in Table 5.
Table 5. Positive and negative samples for homophily and non-homophily methods.
@@
Homophily ratio is calculated by Eq. 1. Note that BZR and COX2 do not have any node label, and thus it is not able to calculate their homophily ratios.
C. Further Descriptions of Datasets
We conduct experiments on ten benchmark datasets. We summary these datasets in Table 6.
  • Wisconsin [^3] https://github.com/bingzhewei/geom-gcn/tree/master/new_data/wisconsin (Pei et al., 2020) is a network of 251 nodes, where each node stands for a webpage, and 199 edges signify the hyperlinks connecting these pages. The features of the nodes are derived from a bag-of-words representation of the webpages. These pages are manually classified into five categories: student, project, course, staff, and faculty. The edge homophily ratio is 0.21.
  • Cornell [^4] https://github.com/bingzhewei/geom-gcn/tree/master/new_data/cornell (Pei et al., 2020) is also a webpage network. It comprises 183 nodes, each symbolizing a webpage, and 295 edges, which represent the hyperlinks between these pages. The node features are obtained from a bag-of-words representation of the webpages. These pages are manually sorted into five categories: student, project, course, staff, and faculty. The edge homophily ratio is 0.22.
  • Chameleon [^5] https://github.com/SitaoLuan/ACM-GNN/tree/main/new_data/chameleon (Rozemberczki et al., 2021) is a Wikipedia network, consisting of 2,277 Wikipedia pages. The pages are divided into five categories according to their average monthly traffic. This dataset creates a network of pages with 36,101 connections, and the node features consist of various key nouns extracted from the Wikipedia pages. The edge homophily ratio is 0.23.
  • Squirrel [^6] https://github.com/SitaoLuan/ACM-GNN/tree/main/new_data/squirrel (Rozemberczki et al., 2021) comprises 5,201 Wikipedia web pages of discussing the defined topics. The dataset is also divided into five categories according to their average monthly traffic. This dataset is a page-page network with 217,073 edges, and the node features are based on several informative nouns in the Wikipedia pages. The edge homophily ratio is 0.30.
  • PROTEINS [^7] https://www.chrsmrrs.com/graphkerneldatasets/PROTEINS.zip (Borgwardt et al., 2005) comprises a dataset of protein graphs, reflecting various characteristics such as amino acid sequences, conformations, structures, and unique features like active sites. In this dataset, each node represents a secondary structure, and each edge signifies a neighboring relationship either within the amino acid sequence or in three-dimensional space. Nodes are classified into three categories, while the graphs themselves are divided into two classes. The edge homophily ratio is 0.66.
  • ENZYMES [^8] http://www.chrsmrrs.com/graphkerneldatasets/ENZYMES.zip (Wang et al., 2022) is a collection of 600 enzymes, sourced from the BRENDA enzyme database. The enzymes are divided into 6 different classes, following their top-level EC enzyme classification. The edge homophily ratio is 0.67.
  • Citeseer [^9] https://nrvis.com/download/data/labeled/citeseer.zip (Sen et al., 2008) contains 3,312 scientific papers, divided into six different categories. The dataset includes a citation network with 4,732 edges. Each paper is represented by a binary word vector, indicating the presence or absence of each word from a dictionary comprising 3,703 unique terms. The edge homophily ratio is 0.74.
  • Cora [^10] https://relational.fit.cvut.cz/dataset/CORA (McCallum et al., 2000) includes 2,708 scientific papers, divided into seven distinct categories. The dataset features a citation network with 5,429 edges. Each paper is represented by a binary word vector, indicating whether each of the 1,433 unique words from the dictionary is present or absent. The edge homophily ratio is 0.81.
  • BZR [^11] https://www.chrsmrrs.com/graphkerneldatasets/BZR.zip (Rossi and Ahmed, 2015) comprises a dataset of 405 ligands linked to the benzodiazepine receptor, represented as individual graph structures. These ligands are divided into 2 distinct categories.
  • COX2121212https://www.chrsmrrs.com/graphkerneldatasets/COX2.zip (Rossi and Ahmed, 2015) consists of a dataset of 467 molecular structures representing cyclooxygenase-2 inhibitors. In this dataset, each node corresponds to an atom and each edge denotes a chemical bond—single, double, triple, or aromatic—between atoms. The molecules are classified into two categories.
@@ C. Further Descriptions of Datasets
1:Pre-trained graph encoder with parameters $\Theta_{0}$,
2:a set of downstream tasks $\mathcal{T}=\{t_{1},\ldots,t_{n}\}$.
3:Optimized parameters $\{\phi_{t_{1}},\ldots,\phi_{t_{n}}\}$ of $n$ condition-nets
4:for $i\leftarrow 1$ to $n$ do
5:    /* Encoding graphs via pre-trained graph encoder */
6:    for each graph $G=(V,E,\mathbf{X})$ in task $t_{i}$ do
7:         $\mathbf{H}\leftarrow\textsc{GraphEncoder}(G;\Theta_{0})$
8:         $\mathbf{h}_{v}\leftarrow\mathbf{H}[v]$, where $v$ is a node in $G$
9:    $\phi_{i}\leftarrow$ initialization
10:    while not converged do
11:         for each node $v\in V$ in task $t_{i}$ do
12:             /* Subgraph sampling and readout Eq. (7) */
13:             Sample $v$’s $k$-hop subgraph $S_{v}$
14:             $\mathbf{s}_{v}\leftarrow\textsc{Average}(\{\mathbf{h}_{u}\cdot\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{v}):u\in V(S_{v})\})$
15:             /* Generate pattern-based prompts by Eq. (8) */
16:             $\mathbf{p}_{t_{i},v}\leftarrow\textsc{CondNet}(\mathbf{s}_{v};\phi_{t_{i}})$
17:             /* Prompt modification by Eq. (9) */
18:             $\tilde{\mathbf{h}}_{t_{i},v}\leftarrow\mathbf{p}_{t_{i},v}\odot\mathbf{h}_{v}$
19:         $\mathbf{h}_{t_{i},G}=\textsc{Average}(\tilde{\mathbf{h}}_{t_{i},v}:v\in\mathcal{V})$
20:         /* Update prototypical subgraphs */
21:         for each class $c$ in task $t_{i}$ do
@@
22:             ${\mathbf{\bar{h}}}_{t_{i},c}\leftarrow\textsc{Average}(\mathbf{\tilde{h}}_{t_{i},x}$: instance $x$ belongs to class $c$)
23:         /* Optimizing the parameters in condition-net */
24:         Calculate $\mathcal{L}_{\text{down}}(\phi_{i})$ by Eq. (10)
25:         Update $\phi_{i}$ by backpropagating $\mathcal{L}_{\text{down}}(\phi_{t_{i}})$
26:return $\{\phi_{t_{1}},\ldots,\phi_{t_{n}}\}$
D. Further Descriptions of Baselines
In this section, we present more details for the baselines used in our experiments.
(1) End-to-end Graph Neural Networks
  • GCN (Kipf and Welling, 2017): GCN utilizes a mean-pooling strategy for neighborhood aggregation to integrate information from neighboring nodes.
  • GAT (Veličković et al., 2018): GAT also leverages neighborhood aggregation for end-to-end node representation learning, uniquely assigns varying attention weights to different neighbors, thereby adjusting their impact on the aggregation process.
  • H2GCN (Zhu et al., 2020b): H2GCN improves node classification by separating ego- and neighbor-embeddings, using higher-order neighborhoods, and combining intermediate representations. These designs help it perform well on both homophilous and heterophilous graphs.
  • FAGCN (Bo et al., 2021): FAGCN improves node representation by adaptively combining low- and high-frequency signals using a self-gating mechanism, making it effective for different network types and reducing over-smoothing.
(2) Graph Pre-training Models
  • DGI (Veličković et al., 2018): DGI operates as a self-supervised pre-training methodology tailored for homogeneous graphs. It is predicated on the maximization of mutual information (MI), aiming to enhance the estimated MI between locally augmented instances and their global counterparts.
  • GraphCL (You et al., 2020): GraphCL leverages a variety of graph augmentations for self-supervised learning, tapping into the intrinsic structural patterns of graphs. The overarching goal is to amplify the concordance between different augmentations throughout graph pre-training.
  • DSSL (Xiao et al., 2022): DSLL uses latent variable modeling to decouple semantics in neighborhoods, avoiding augmentations and optimizing with variational inference to capture local and global information, enhancing node representation learning.
  • GraphACL (Xiao et al., 2023): GraphACL considers each node from two perspectives: identity representation and context representation. The model trains the former by predicting the context representation of one-hop neighbors using an asymmetric predictor, and then reconstructs the same latter of the central node by enforcing identity representations from two-hop neighbors.
(3) Graph Prompt Models
  • GPPT (Sun et al., 2022): GPPT utilizes a GNN model pre-trained via a link prediction task which is a strong homophily method. The downstream prompt module is designed specifically for node classification, aligning it with the pre-training link prediction task.
  • GraphPrompt (Liu et al., 2023a): GraphPrompt employs subgraph similarity calculations as a unified template to bridge the gap between pre-training and downstream tasks, including node and graph classification. A learnable prompt is fine-tuned during downstream adaptation to incorporate task-specific knowledge.
  • GraphPrompt+ (Yu et al., 2024c): GraphPrompt+ builds on GraphPrompt by introducing a series of prompt vectors within each layer of the pre-trained graph encoder. This technique utilizes hierarchical information from multiple layers, beyond just the readout layer.
E. Implementation Details of Approaches
General settings Optimizer. For all experiments, we use the Adam optimizer.
Environment. The environment in which we run experiments is:
  • Linux version: 5.15.0-78-generic
  • Operating system: Ubuntu 18.04.5 LTS
  • CPU information: Intel(R) Xeon(R) Platinum 8352V
  • GPU information: GeForce RTX 4090 (24 GB)
Figure 5. Impacts of different shots on node and graph classification.
@@ E. Implementation Details of Approaches
Details of baselines. We use the official code provided for all open-source baselines. Each model is tuned according to the settings recommended in their respective publications to ensure optimal performance. We use early stopping strategy for training and set patience to 50 steps. The number of training epochs is set to 2,000.
@@
  • For the baseline GCN (Kipf and Welling, 2017), we employ a 3-layer architecture on Wisconsin, Squirrel, Chameleon, Cornell datasets and 2-layer architecture on Cora, Citeseer, ENZYMES, PROTEINS, COX2, BZR datasets. Hidden dimensions is 256.
@@
  • For FAGCN (Bo et al., 2021), we employ a 2-layer architecture. The hyper-parameter setting is: eps = 0.3, dropout = 0.5, hidden = 256. we use relu as activation function.
  • For DGI (Veličković et al., 2018), we utilize a 1-layer GCN as the base model and set the hidden dimensions to 256. Additionally, we employ prelu as the activation function.
  • For GraphCL (You et al., 2020), a 1-layer GCN is also employed as its base model, with the hidden dimensions set to 256. Specifically, we select edge dropping as the augmentations, with a default augmentation ratio of 0.2.
  • For DSSL (Xiao et al., 2022), the hidden dimension search space is in {64, 256, 2048}. We report the best performance on PROTEINS and ENZYMES with hidden size of 64 , Cora and Citeseer with 2048, and the rest datasets with 256. We keep the other hyper-parameters the same as in the original demonstrations in their Github repository.
  • For GraphACL (Xiao et al., 2023),the hidden dimension search space is in {64, 256, 1024, 2048}. We report the best performance on PROTEINS and ENZYMES with hidden size of 64 , Cora and Citeseer with 2048, and the rest datasets with 256. We keep the other hyper-parameters the same as in the original demonstrations in their Github repository.
  • For GPPT (Sun et al., 2022), we utilize a 2-layer GraphSAGE as its base model, setting the hidden dimensions to 256. For base GraphSAGE, we also employ a mean aggregator.
  • For GraphPrompt (Liu et al., 2023a), we employ a 3-layer architecture on Wisconsin, Squirrel, Chameleon, Cornell datasets and 2-layer architecture on Cora, Citeseer, ENZYMES, PROTEINS, COX2, BZR datasets. Hidden dimensions are set to 256.
  • For GraphPrompt+ (Yu et al., 2024c), we employ a 2-layer GCN on Cora, Citeseer, ENZYMES, PROTEINS, COX2, BZR datasets and 3-layer GCN on the rest datasets. Hidden dimensions are set to 256.
Details of ProNoG. For our proposed ProNoG, we utilize a 2-layer FAGCN architecture as backbone for pre-training task with graph contrastive methods for Wisconsin, Squirrel, Chameleon, Cornell. Especially, we implement edge-dropping on sub-graph level for Wisconsin, Squirrel, Chameleon, Cornell. Hidden dimensions are set to 256. For Cora, Citeseer, BZR, COX2, we employ 1-layer GCN as base model for pre-training task. Hidden dimensions are set to 256. For PROTEINS, we employ 1-layer GCN on link prediction task for pre-training. Hidden dimensions is set to 64. For ENZYMES, we implement DSSL for pretraining. Hidden dimensions is set to 64. All experiments are undertaken with the seed of 39. Especially, we found that on Chameleon, Squirrel, keeping the original node features as input without normalization performs the best, while for others, normalization of node features remains routine. Except for DSSL, we use cosine-similarity loss on node level as loss function.
@@
ProNoG 1,024 1,024 1,024 1,024
F. Impact of Shots
We vary the number of shots and conduct node classification on Cornell and Citeseer, and graph classification tasks on Wisconsin, Cornell, BZR and COX2. The results are illustrated in Fig. 5, and we observe same patterns as shown in node classification tasks on other datasets.
G. Parameters efficiency
We evaluate the parameter efficiency of ProNoG compared to other notable methods. Specifically, we evaluate the number of parameters that need to be updated or tuned during the downstream adaptation phase, and present the results in Table 7. For GCN and FAGCN, since these models are trained end-to-end, all model weights must be updated, leading to the lowest parameter efficiency. In contrast, for GraphCL and xxx, only the downstream classifier is updated, while the pre-trained model weights remain unchanged, significantly reducing the number of parameters that require tuning. Prompt-based methods GraphPrompt, GraphPrompt+, and ProNoG are the most parameter-efficient, as prompts or condition-net are lightweight and contain fewer parameters than typical classifiers like fully connected layers. Note that the reported number of parameters for ProNoG are based on $d=2$, given that ProNoG still performs competitively with such hyperparameter setting. Although our conditional prompt design requires to update more parameters than GraphPrompt and GraphPrompt+ during downstream adaptation, the increase is minor compared to updating the entire classifier or model weights, and thus does not pose a major issue.
H. Hyperparameter Analysis
In our experiment, we use a 2-layer MLP with a bottleneck structure as the condition-net. We evaluate the impact of the hidden dimension of the condition-net $m$ and report the corresponding performance in Fig. 6. We observe that for both node and graph classification, as $m$ increases from 2, the performance generally first decreases because a larger $m$ introduces more learnable parameters, which may lead to worse performance in few-shot settings. However, after reaching a trough, accuracy starts to gradually increase as $m$ grows further, since higher dimensions increase model capacity, until reaching a peak. Then the performance further declines as $m$ improves, given more learnable parameters. Note that the overall variation in performance is generally small, and the peak is generally at $m=2$ or $m=64$. In our experiment, we set $m=64$ in our experiments.
Figure 6. Sensitivity study of $m$.
I. Further Comparison with Other Baselines
We conduct further one-shot node classification on Wisconsin and compare with other ”pre-training, prompting” baselines. The results are shown in Table 8. ProNoG consistently outperforms these methods, further demonstrating the effectiveness of ProNoG
Table 8. Accuracy evaluation on one-shot node classification.
Methods Wisconsin
Self-pro (Gong et al., 2024) 29.81 $\pm$ 10.33
PSP (Ge et al., 2024) 33.71 $\pm$ 7.16
Prodigy (Huang et al., 2024) 25.34 $\pm$ 0.63
All-in-one (Sun et al., 2023a) 23.59 $\pm$ 0.10
ProNoG 44.72 $\pm$ 11.93
J. Efficiency Analysis
We compare the prompt-tuning time of ProNoG and a competitive baseline GraphPrompt (Liu et al., 2023a) and show the results in Table 9. Empirically, our prompt tuning takes only marginally longer than the strong baseline GraphPrompt, as shown below (in seconds), while significantly outperforming GraphPrompt in terms of performance. Therefore, this trade-off is acceptable.
Table 9. Prompt-tuning times (seconds) on node classification.
Methods Cora
GraphPrompt 1.003
ProNoG 1.513
K. Performance on Large-scale Dataset
We evaluate ProNoG under the same setting as introduced in Section 6.1 on a large-scale dataset Amazon (Zeng et al., 2019), which consists of 1,598,960 nodes, 132,169,734 edges, and 107 classes. The results are shown in Table 10. We observe that ProNoG outperforms competitive baselines, showing the effectiveness of ProNoG on large-scale dataset.
Table 10. Accuracy evaluation on one-shot node classification.
Methods Amazon
GraphPrompt 3.91
ProNoG 4.56

### v6

@@ A. Algorithm
We detail the main steps for conditional prompt generation and tuning in Algorithm 1. In brief, we iterate through each downstream task to learn the corresponding prompt vectors. In lines 3–5, we compute the embedding for each node using the pre-trained graph encoder, with the pre-trained weights $\Theta_{0}$ frozen throughout the adaptation process. In lines 6–22, we optimize the condition-net. Specifically, we perform similarity-weighted readout (lines 9–11), generate prompts (lines 12–13), modify nodes’ embeddings using these prompts (lines 14–15), calculate the embedding of the corresponding graph (line 16), and update the embeddings for the prototypical nodes/graphs based on the few-shot labeled data provided in the task (lines 17–19).
@@ B. Homophily and Non-Homophily Methods
We provide further details about the set of positive samples $\mathcal{A}$ and negative samples $\mathcal{B}$ for various homophily and non-homophily methods in Table 5.
Table 5. Positive and negative samples for homophily and non-homophily contrastive methods.
@@
Homophily ratios are calculated by Eq. (1). Note that BZR and COX2 do not have any node label, and thus no homophily ratios can be calculated.
C. Further Experimental Details
Datasets. We summarize the statistics of the ten datasets used in our experiments in Table 6.
@@ C. Further Experimental Details
1:Pre-trained graph encoder with parameters $\Theta_{0}$, a set of downstream tasks $\mathcal{T}=\{t_{1},\ldots,t_{n}\}$.
2:Optimized parameters $\{\phi_{t_{1}},\ldots,\phi_{t_{n}}\}$ of $n$ condition-nets
3:for $i\leftarrow 1$ to $n$ do
4:    /* Encoding graphs via pre-trained graph encoder */
5:    for each graph $G=(V,E,\mathbf{X})$ in task $t_{i}$ do
6:         $\mathbf{H}\leftarrow\textsc{GraphEncoder}(G;\Theta_{0})$
7:         $\mathbf{h}_{v}\leftarrow\mathbf{H}[v]$, where $v$ is a node in $G$
8:    $\phi_{i}\leftarrow$ initialization
9:    while not converged do
10:         for each node $v\in V$ in task $t_{i}$ do
11:             /* Subgraph sampling and readout by Eq. (7) */
12:             Sample $v$’s $k$-hop subgraph $S_{v}$
13:             $\mathbf{s}_{v}\leftarrow\textsc{Average}(\{\mathbf{h}_{u}\cdot\mathtt{sim}(\mathbf{h}_{u},\mathbf{h}_{v}):u\in V(S_{v})\})$
14:             /* Generate pattern-based prompts by Eq. (8) */
15:             $\mathbf{p}_{t_{i},v}\leftarrow\textsc{CondNet}(\mathbf{s}_{v};\phi_{t_{i}})$
16:             /* Prompt modification by Eq. (9) */
17:             $\tilde{\mathbf{h}}_{t_{i},v}\leftarrow\mathbf{p}_{t_{i},v}\odot\mathbf{h}_{v}$
18:         $\mathbf{h}_{t_{i},G}=\textsc{Average}(\tilde{\mathbf{h}}_{t_{i},v}:v\in\mathcal{V})$
19:         /* Update prototypical subgraphs */
20:         for each class $c$ in task $t_{i}$ do
@@
21:             ${\mathbf{\bar{h}}}_{t_{i},c}\leftarrow\textsc{Average}(\mathbf{\tilde{h}}_{t_{i},x}$: instance $x$ belongs to class $c$)
22:         /* Optimizing the parameters in condition-net */
23:         Calculate $\mathcal{L}_{\text{down}}(\phi_{i})$ by Eq. (10)
24:         Update $\phi_{i}$ by backpropagating $\mathcal{L}_{\text{down}}(\phi_{t_{i}})$
25:return $\{\phi_{t_{1}},\ldots,\phi_{t_{n}}\}$
@@
Details of baselines. We use the authors’ code for all baselines, if available. To ensure a fair comparison, each model is tuned while referencing the settings recommended in their respective publications. We use early stopping strategy in training and set the patience value to 50 steps. The number of training epochs is set to 2,000.
@@
  • For GCN (Kipf and Welling, 2017), we employ a 3-layer architecture on Wisconsin, Squirrel, Chameleon, Cornell datasets and a 2-layer architecture on Cora, Citeseer, ENZYMES, PROTEINS, COX2, BZR datasets. The hidden dimension is set to 256.
@@
  • For FAGCN (Bo et al., 2021), we employ a 2-layer architecture. The hyper-parameter settings are: eps = 0.3, dropout = 0.5, hidden = 256. We use relu as the activation function.
  • For DGI (Veličković et al., 2018), we utilize a 1-layer GCN as the base model and set the hidden dimension to 256. Additionally, we employ prelu as the activation function.
  • For GraphCL (You et al., 2020), a 1-layer GCN is also employed as its base model, with the hidden dimension set to 256. Specifically, we select edge dropping as the augmentation, with a default augmentation ratio of 0.2.
  • For DSSL (Xiao et al., 2022), we search the hidden dimension in {64, 256, 2048}. We report the best performance on PROTEINS and ENZYMES with a hidden size of 64, Cora and Citeseer with 2048, and the rest datasets with 256.
  • For GraphACL (Xiao et al., 2023), we search the hidden dimension in {64, 256, 1024, 2048}. We report the best performance on PROTEINS and ENZYMES with a hidden size of 64, Cora and Citeseer with 2048, and the rest datasets with 256.
  • For GPPT (Sun et al., 2022), we utilize a 2-layer GraphSAGE as its base model, setting the hidden dimensions to 256. For the GraphSAGE backbone, we employ a mean aggregator.
  • For GraphPrompt (Liu et al., 2023a), we employ a 3-layer architecture on Wisconsin, Squirrel, Chameleon, and Cornell, and a 2-layer architecture on the rest. Hidden dimensions are set to 256. We use link prediction as the pre-training task.
  • For GraphPrompt+ (Yu et al., 2024c), we employ a 2-layer GCN on Cora, Citeseer, ENZYMES, PROTEINS, COX2, and BZ, and a 3-layer GCN on the rest. Hidden dimensions are set to 256. We use link prediction as the pre-training task.
Details of ProNoG. For our proposed ProNoG, we utilize a 2-layer FAGCN architecture as the backbone for pre-training on the non-homophilic graphs, namely, Wisconsin, Squirrel, Chameleon, and Cornell, with edge-dropping implemented on the subgraph level and hidden dimensions set to 256. For the remaining more homophilic graphs, we employ a 1-layer GCN for pre-training, with the hidden dimensions set to 256, except for PROTEINS and ENZYMES, which use hidden dimensions of 64. We adopt a non-homophily pre-training task GraphCL (You et al., 2020) for all datasets except for PROTEINS and ENZYMES. Specifically, GraphCL does not work well for ProNoG on the two datasets, and instead we use link prediction (Liu et al., 2023a) and DSSL (Xiao et al., 2022), respectively. Note that the non-homophily task GraphCL still works well on most datasets including all of the non-homophilic graphs. For the condition-net, we set the hidden dimension to 64 for all datasets. All experiments are conducted with a random seed of 39.
@@
ProNoG 32,768 32,768 32,768 32,768
D. Parameter Efficiency
We evaluate the parameter efficiency of ProNoG compared to other notable methods. Specifically, we evaluate the number of parameters that need to be updated or tuned during the downstream adaptation phase, and present the results in Table 7. For GCN and FAGCN, since these models are trained end-to-end, all model weights must be updated, leading to the least parameter efficiency. In contrast, for GraphCL and GraphACL, only the downstream classifier is fine-tuned, while the pre-trained model weights are frozen, significantly reducing the number of tunable parameters in the downstream phase. Prompt-based methods GraphPrompt and GraphPrompt+ are the most parameter-efficient, as prompts are lightweight and typically contain fewer parameters than the downstream classifier. Although our conditional prompt design requires updating more tunable parameters than GraphPrompt and GraphPrompt+ during downstream adaptation, the increase is minor compared to updating the entire model weights, and thus does not pose a major issue.
Figure 5. Impact of hidden dimension $m$ in the condition-net.
E. Hyperparameter Analysis
In our experiments, we use a 2-layer MLP with a bottleneck structure as the condition-net. We evaluate the impact of the hidden dimension $m$ of the condition-net, and report the corresponding performance in Fig. 5. We observe that for both node and graph classification, $m=64$ generally yields optimal performance, which we have adopted in all other experiments. Specifically, smaller values of $m$ may lack sufficient model capacity, while larger $m$ may introduce too many learnable parameters, leading to overfitting in few-shot settings.